how many quarts of 6% butterfatmilk and 3% of butterfat milk should be mixed to yield 75 quarts of 5%butterfat milk
Answer by mananth(16946)

............... percent ----------------quantity
butterfat milk 6----------------x quarts
butterfat milk 3----------------75-x quarts
Total 5.00% ---------------- 75
6x+3(75 -x)=5*75
6x+225-3x=375
6x-3x=375-225
3x=150
/3
x=50 quarts 6.00%